the intensity ratio of two sound waves is 1:9 the ratio of their amplitude is

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
Intensity proportional to a²
I1/ I2 = a²1/a²2
 
1/9 = (a1/a2)²
 
Hence
 
a1/a2 = 1/3
 
 
Regards
Arun (askIITians forum expert)